OpenStudy (anonymous):

how to multiply (2x^2-5x+6)(3x=2) please explain

sam (.sam.):

Can you rewrite that?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I need to find the product of (2x^2-5x+6)(3x-2)

mathslover (mathslover):

@userhaha , you can also use the "equation" tag given just left to the "draw" and "attach file" tags to write the equations in a beautiful form and much understandable way , it is. \(\large(2x^2 - 5x + 6)(3x-2) \) {Just like I did :) }

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

can you explain to me how to find the product?

mathslover (mathslover):

Yes.

mathslover (mathslover):

Do you know distributive property?

OpenStudy (anonymous):

yes

mathslover (mathslover):

Good. Can you say what does it state?

OpenStudy (anonymous):

isn't it when you break up the problem into parts and then combine them using like terms?

mathslover (mathslover):

I can write that in equation form as : \(a(b+c) = ab + ac\) or \((a+b)(c+d) = ac + ad + bc + bd\)

mathslover (mathslover):

Now, we have : \((2x^2 - 5x +6)(3x-2)\) Here we are also going to use distributive property to find the product. Can you say me , how will you expand this using distributive property ?

OpenStudy (anonymous):

\[2x^2(3x)+2x^2(-2)+-5x(3x)+-5x(-2)+6(3x)+6(-2)\]

OpenStudy (anonymous):

Is that correct?

mathslover (mathslover):

Very good. That's correct. Now just solve it! Can you?

OpenStudy (anonymous):

That's where I was thinking I was doing it wrong. I guess I've been doing my homework correctly then. I'll solve it right now

OpenStudy (anonymous):

\[6x^3-19x^2+28x-12\]

mathslover (mathslover):

Openstudy is a right place to consult doubts and problems encountered while doing questions. Good to see you're using OS in a right manner.

mathslover (mathslover):

\(\color{gold}{\mathsf{Excellent}} \ddot \smile \) You're right.

OpenStudy (anonymous):

Thank you

mathslover (mathslover):

You're welcome . I suggest you to also have a look here : http://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/ To begin your journey in OS, the above guide is the best to start with. It will guide you about the basic rules and options available like drawing, attaching files, writing equations etc. best of luck